Session: Feb 1999 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Thu Feb 25 17:16:07 MET 1999) Minor failures
a couple of scans lost due to problems with the field system

Torun (K.Borkowski at Thu Feb 25 18:29:09 MET 1999) Minor failures
One hour (11:50 to 12:50) observations off source due to malfunction of the elevation drive.

Noto (C.Trigilio at Fri Feb 26 11:26:46 MET 1999) Success

Effelsberg (dgraham at Fri Feb 26 12:49:48 MET 1999) Minor failures
Puzzled by abscence of lines on autocorrelation, possible problem with dopsrc in schedule ? Otherwise OK

Onsala (A. Polatidis at Fri Feb 26 14:11:18 MET 1999) Success

Westerbork (Tony Foley at Fri Feb 26 22:36:56 MET 1999) Minor failures
Lots of EMI around 1625MHz. Lost 0900-0912UT as the telescope control computer crashed (a memory leak caused a new version of the receiver control program to grow continuously).

Jodrell (Alastair Gunn at Tue Mar 2 05:39:02 MET 1999) Minor failures
Missed start of scan at 09:42 due to long slew, on source at 09:53 UT. Elevation servo fault probably lost five minutes around 13:15 UT. No other known problems.

Calibration (Jean\-Francois Desmurs at Wed Mar 3 18:33:05 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: ey004a.antab. Read the comments in the file to get further information.